OpenStudy (austinl):

Solve the given differential equation. \(y\prime = \dfrac{(3x^2-1)}{3+2y)}\) Where would I begin?

OpenStudy (anonymous):

As simple as this:\[y'=\frac{ dy }{ dx }=\frac{ 3x^2-1 }{ 3+2y } \rightarrow (3+2y)dy=(3x^2-1)dx\]You can integrate

OpenStudy (austinl):

Oh my goodness. That makes gobs more sense.

OpenStudy (austinl):

Ok, so now I have: \(y^2+3y+C=x^3-x+C\) Now, I haven't a clue what to do next, combine like terms and solve for y?

OpenStudy (austinl):

\(\large{y=\pm \dfrac{-(\sqrt{4x^3-4x+9}+3)}{2}}\) That makes absolutely zero sense.

OpenStudy (austinl):

+C on the end of course....

OpenStudy (loser66):

lol!!! Actually, you don't need to find y out, just let it there as y ^2 +3y -x^3 +x = C that's it

OpenStudy (austinl):

Many typos there. \(y^2+3y-x^3+x=C~,~y\ne\frac{-3}{2}\)

OpenStudy (loser66):

it's ok, looks good to me

OpenStudy (austinl):

Woohoo, very nice!

OpenStudy (loser66):

hahahaha... glad to see "Woohoo..." wudwud... You are gud

OpenStudy (anonymous):

If you leave the solution like: \[3y+y^2-x^3+x=C\]is perfectly valid